If consumption is 1800, GDP is 4300, government purchases are 1000, imports are 700, and investments is 1200, then exports are 300?
Answer #1

GDP = Consumption + government expenditure + investment + (Exports - imports)

4300 = 1800 + 1000 + 1200 + (300 - 700) = 4000 - 400 = 3600.

4300 - 3600 = 700.

The above statement is false. For the GDP to be 4300 the exports should be 300 + 700 = 1000.